OpenStudy (anonymous):

If the level of a lake changes by -2 inches each year for 6 years, what will the change in lake level be at the end of those 6 years?

OpenStudy (mathmate):

-2 first year -2-2=-4 second year -2-2-2=-6 third year -2-2-2-2=-8 fourth year -2-2-2-2-2=-10 fifth year what is the sixth year?

OpenStudy (anonymous):

-2 -2 -2 -2 -2 -2 -2 =-12 sixth year

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Sounds good to me! :)

OpenStudy (anonymous):

@mathmate ty

OpenStudy (mathmate):

yw!